OpenStudy (anonymous):

Which of the following systems of equations has infinitely many solutions? 4x − 7y = 1; 14y = 8x − 3 8x − 3 = 6y; 12y − 6 = 2x 4x − 5y = 13; 10y = 8x − 26 5y = -x − 5; -2x + 15y = 10

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

The others are either 1 line or crossing

OpenStudy (anonymous):

@Ashley1997 Isn't it the third one? You get infinitely many solutions when it's identity. 4x - 5y = 13 -8x + 10y = -26 8x - 10y = 26 0 = 0

OpenStudy (mertsj):

The first one are two parallel lines.

OpenStudy (anonymous):

That would be no solution. It asks for infinitely many solutions which is an identity. 0 = 0. Not 0 = 6 or any of those.